Grade 12Physical Chemistry

Density of a gas is found to be 5.46 g/dm3 at 27 °C at 2 bar pressure. What will be its density at STP?

sudhanshu
12 Years agoGrade 12

4 Answers

Gaurav
11 Years ago
Hello Student
P1V1/T1 = P2V2/T2
2V1/300 = 1V2/273
(273/150)V1 = V2
1.82V1 = V2
Density at STP = mass/volume = m/1.82V1 = (5.46/1.82) g/dm3 = 3 g/dm3
Vishnu kumar dev
7 Years ago
D=pM/RT
5.46=2×M×R×T+27...........(1)
D=1×M×R×273k..............(2)
Upon Dividing both equations we will get D=3dm*3
Mohammad shoeb
7 Years ago
d1/d2 = P1/P× T2/T1........................by relation d=PM/RT
5.46/d= 2/1 × 273/300
d2 = 3dm3............👍👍👍👍👍👍👍
 
Rishi Sharma
6 Years ago
Dear Student,
Please find below the solution to your problem.

P1V1/T1 = P2V2/T2
2V1/300 = 1V2/273
(273/150)V1 = V2
1.82V1 = V2
Density at STP = mass/volume
= m/1.82V1
= (5.46/1.82) g/dm^3
= 3 g/dm^3

Thanks and Regard